Subject: APO 3.5m users committee minutes, Nov 2014 meeting

From: Suzanne Hawley

Submitted: Sun, 7 Dec 2014 23:08:04 -0800 (PST)

Message number: 1348 (previous: 1347, next: 1349 up: Index)

 		APO 3.5-m Users Committee Phonecon, 11/17/14

Attending:  Suzanne Hawley, Mark Klaene, Bill Ketzeback, Joe Huehnerhoff, 
Scott Anderson, Rachel Kuzio de Naray, Joanne Hughes Clark, John 
Wisniewski, Jon Holtzman, David Brown (for Michael Strauss), Mark 
Hammergren, Jeff Bary, John Bally, Rene Walterbos, and Bruce Gillespie


User feedback and comments from institutional representatives:

o Washington - Scott had no issues or concerns.

o Georgia State - Rachel had no issues or concerns.

o Seattle - Joanne reported that there was a pointing/software issue on 
one night that was later resolved.  SPIcam worked well, and the seeing was 

o Oklahoma - John had no issues or concerns.

o Princeton - David said that the PU users have no reported issues.

o NMSU - Jon had no issues or concerns.

o Chicago/Adler - Mark mentioned problems with the new Chebyshev tracking 
routines; these are being worked on with Russell Owen.

o NAPG - Jeff said the users are happy.

o Virginia - No report.

o Colorado - John had no user concerns.

o Johns Hopkins - No report.


Discussion of telescope/instruments report:

Mark highlighted the report below. There were two occurrences of issues 
with the azimuthal motion of the telescope, which are now understood. 
Also, TripleSpec twice produced corrupted images, and then corrected 
itself in later image files. Staff at Virginia are looking into it.  Joe 
reported that the new imager, ARCTIC, is currently pumped down, powered 
on, and he is starting the characterization of the CCD.  Bill reported 
that ARCSAT has been behaving well, although there have been a couple of 
instances of software glitches.  It seems that most ARCSAT users have been 
happy with the shared-risk observing program.


              3.5-m Telescope and Instruments Highlights
                        10/22/14 through 11/14/14

0) Overview

Weather for the end of October and the first half of November has started 
to cool off with several nights with high humidity.  We have yet to see 
our first snow fall of the year.

1) Telescope

We had two instances of telescope runaways.  We believe we understand each 
instance and have taken corrective measures in software.  Otherwise the 
telescope has been behaving well with the new TCC software.

2) Instruments

DIS has been operating normally.

Agile is operational.

TripleSpec has had two instances of corrupted images.  The first instance 
the problem corrected itself before further troubleshooting could be done. 
The second is being investigated.  Mike Skrutskie is on site and we are 
consulting with him.

SPIcam was operational during this period.

Echelle inter-order light is holding with no additional issues.

NIC-FPS continues to operate in shared-risk observing mode due to 
occasional image corruption issue.  NFS less than 9 is operational.

GIFS is operational.


The instrument development is proceeding on schedule.  The dewar has a 
loaner CCD from STA and has been integrated with the readout electronics 
by Bob Leach.  It is at UW for further testing and development.  The 
optics order will be placed within the next few weeks.  We hope to start 
fabricating the remainder of the instrument after the first of the year.


Report on Board meeting:

Suzanne said that the ARC Board of Governors meeting went well, and thanks 
for all the slides that showed scientific results from the past year. 
Notable in the past year was the unusually bad weather in September. The 
overall downtime rate of 0.6% is similar to the last few years and is 
again an impressive achievement and tribute to the 3.5m staff.  The new 
TCC and ARCTIC development were also highlighted.  It was pointed out that 
with the new GSU partnership and existing leases, we are now financially 
solvent through 2016, and the plan is that extra income will be invested 
in infrastructure and instrumentation upgrades.  Rene added that the Board 
was pleased with the stability of the telescope partnership and its 
operations, as well as the new developments either underway or being 

Mark led a discussion at the Board meeting on the pending situation at 
Sunspot with the National Solar Observatory.  Although the date may slip, 
it is currently believed that the NSF will relieve NSO of responsibility 
for the Sunspot site in 2017.  The biggest impact to APO will be access to 
water.  Other concerns entail use of the aluminizing chamber, housing for 
staff and visitors, and the occasional use of heavy equipment.  All these 
are at risk, and we are discussing with NMSU a proposal to take over 
limited operations at Sunspot of the critical facilities and services we 
need.  This said, it is apparently not completely clear to us what are the 
intentions of the NSF when NSO leaves.


GIFS update:

Mark announced that GIFS is up and running, and is being kept cold.  The 
instrument is available for remote use on the telescope, and there is 
on-line documentation.  Engineers at GSFC are planning to bring a 
photon-counting detector for engineering in 2015.  If interested, contact 
John Wisniewski at Oklahoma or Michael McElwain at GSFC.  It is hoped that 
the new detector will be at APO in the second quarter.  Suzanne added that 
she hopes that ARC users will start to use GIFS as a real facility 
instrument, and potential users should try it out to learn its 


APOGEE fiber link to 3.5m (action item):

Suzanne said she is still polling the community to see how much interest 
there is for a fiber run from the 3.5-m telescope to the H-band APOGEE 
spectrograph at the Sloan telescope.  Based on discussions at the Board 
meeting, it appears that SDSS is OK with the concept of the instrument 
being scheduled for use with the 3.5-m telescope perhaps 2-3 nights/month. 
The Users Committee members relayed the levels of interest for this 
capability at their respective institutions:

   o Scott (UW) - mild interest
   o Joanne (Seattle) - mild interest
   o Rachel (GSU) - none to lukewarm
   o JohnW (Okla) - mild
   o David (Princeton) - no update from last month
   o Jon (NMSU) - Strong interest, especially potential for Gaia follow-up
   o Jeff (NAPG) - young stars, mild interest
   o John (Colorado) - Strong interest, two specific science cases
   o Don York (Chicago) - Strong interest

Suzanne added that we could consider laying fibers between all the APO 
telescopes, which would provide additional opportunities.  Suzanne asked 
for two people to work with Nick MacDonald at UW to develop a technical 
and scientific case for the fiber project, either as internally funded or 
with new funding from external sources.  John Bally and Jon Holtzman 
agreed to be on the panel.  Bruce also volunteered, and Don York has 
expressed interest.  Any other interested users are welcome to join, 
please send email to Jon H. and Suzanne.


New optical spectrograph design study:

Suzanne said that she talked with the ARC Board about the need to replace 
DIS in the next few years.  Before developing a concrete plan for a new 
spectrograph, we first need to have a group to study what characteristics 
and science a new spectrograph would address, which would lead to 
instrumental design requirements.  We need to find a Principal 
Investigator willing to lead the design and construction of the new 
spectrograph which will likely be a $2M project.  This led to an action 

=====> ACTION:  Users Committee members to poll their constituents to 
engage and enlist people who are willing to be involved with the 
specification and development of a new optical spectrograph.



Suzanne informed the Board at the recent meeting that she intends to step 
down as Director of the ARC 3.5-m telescope, likely when her current (4th) 
term ends in December, 2016.  A Director search will probably be initiated 
sometime next year, so there will be some overlap in 2016.  If anyone is 
potentially interested, Suzanne would be happy to discuss the job with 
them even before the formal search process begins.



[open from previous meeting]:

=====> Action: users committee to gather information from users about 
scientific interest in a fiber feed to APOGEE (2.5m) and smaller 
telescopes on site.

 	Status: Commmittee appointed, other users welcome to join, action 
item closed.

[new from this meeting]:

=====> ACTION:  Users Committee members to poll their constituents to find 
people who are willing to be involved with the specification and 
development of a new optical spectrograph.


Next meeting:  The next Users Committee phonecon will be on Monday 15 
December.  The agenda and other materials will be sent to the committee 
members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1348 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to